# A Unified Model for Multi-Task Drone Routing in Post-Disaster Road Assessment

### v1

@@
[…]
A Unified Model for Multi-Task Drone Routing in Post-Disaster Road Assessment
[…]
Post-disaster road assessment (PDRA) is essential for emergency response, enabling rapid evaluation of infrastructure conditions and efficient allocation of resources. Although drones provide a flexible and effective tool for PDRA, routing them in large-scale networks remains challenging. Traditional optimization methods scale poorly and demand domain expertise, while existing deep reinforcement learning (DRL) approaches adopt a single-task paradigm, requiring separate models for each problem variant and lacking adaptability to evolving operational needs. This study proposes a unified model (UM) for drone routing that simultaneously addresses eight PDRA variants. By training a single neural network across multiple problem configurations, UM captures shared structural knowledge while adapting to variant-specific constraints through a modern transformer encoder-decoder architecture. A lightweight adapter mechanism further enables efficient finetuning to unseen attributes without retraining, enhancing deployment flexibility in dynamic disaster scenarios. Extensive experiments demonstrate that the UM reduces training time and parameters by a factor of eight compared with training separate models, while consistently outperforming single-task DRL methods by 6–14% and traditional optimization approaches by 24–82% in terms of solution quality (total collected information value). The model achieves real-time solutions (1–10 seconds) across networks of up to 1,000 nodes, with robustness confirmed through sensitivity analyses. Moreover, finetuning experiments show that unseen attributes can be effectively incorporated with minimal cost while retaining high solution quality. Validation on a real-world transportation network further demonstrates the model’s practical scalability. The proposed UM advances neural combinatorial optimization for time-critical applications, offering a computationally efficient, high-quality, and adaptable solution for drone-based PDRA.
[…]
Post-disaster road assessment, Drone routing, Deep reinforcement learning, Multi-task learning, Lightweight finetuning, Neural combinatorial optimization
@@ 1 Introduction
Disasters, such as earthquakes and hurricanes, pose severe threats to human lives, infrastructure, and economic stability (Avishan et al. 2023). In the immediate aftermath of such events, uncertainties regarding transportation infrastructure disruptions emerge, encompassing the extent of road network damage, operability, and feasibility of short-term restoration. Therefore, rapid assessment of road network damage is critical for effective emergency response, as it enables timely delivery of supplies, evacuation planning, and resource allocation (Zhang et al. 2023). This process, referred to as post-disaster road assessment (PDRA), is hindered by traditional ground-based methods, which are often impeded by impassable roads, hazardous conditions, and delays, making them impractical for time-sensitive disaster scenarios (Zhang et al. 2023, Gong et al. 2025). Unmanned aerial vehicles (or drones) have emerged as a transformative solution for PDRA, offering high mobility, rapid deployment, and the ability to capture high-resolution imagery in hazardous environments (Enayati et al. 2023, Shi et al. 2024, Steenbergen et al. 2025). Equipped with advanced sensors, drones can efficiently assess road damage severity, operability, and repair feasibility, providing critical data to guide relief efforts. Examples include their use in post-typhoon Haiyan (Ezequiel et al. 2015) and the Nepal earthquake (Abuali and Ahmed 2025).
@@
Despite these advantages, solving drone routing for large-scale PDRA remains challenging. Traditional approaches, including exact methods (e.g., branch-and-bound algorithms) (Zhang et al. 2023, Yin et al. 2023, Morandi et al. 2024) and heuristic methods (e.g., simulated annealing) (Oruc and Kara 2018, Glock and Meyer 2020, Adsanver et al. 2025), suffer from significant limitations in real-world large-scale PDRA. These methods exhibit solution times that grow dramatically with problem complexity, making them ill-suited to time-critical disaster response where delays can be life-threatening. Moreover, developing effective solving algorithms requires extensive domain knowledge and substantial time investment in algorithm design. To address these two issues, Gong et al. (2025) proposed a deep reinforcement learning (DRL) approach for PDRA, which learns high-quality routing strategies directly from data without relying on domain knowledge, achieving real-time inference (1–10 seconds) and outperforming commercial solvers in solution quality. In recent years, DRL has been used to learn neural network-based heuristics for solving combinatorial optimization problems, receiving growing research attention for its potential to generate high-quality solutions with minimal human effort (Vinyals et al. 2015, Bello et al. 2016, Kool et al. 2018, Kwon et al. 2020, Luo et al. 2023, Zhou et al. 2024, Liu et al. 2024, Berto et al. 2024). However, this DRL method operates in a single-task paradigm, requiring a separate model for each specific PDRA variant. When problem characteristics change, such as variations in objective functions or constraints, a new model should be trained from scratch, leading to high computational costs and deployment inefficiencies. Furthermore, the existing method lacks adaptability to unseen future attributes in evolving PDRA scenarios.
@@
Inspired by advancements in large language models (LLMs), where a single model can be applied to diverse tasks with remarkable adaptability and deployment efficiency (Achiam et al. 2023, Touvron et al. 2023, Guo et al. 2025, Huang et al. 2025), this study extends the philosophy to neural combinatorial optimization for PDRA. The main contributions are summarized as follows:
@@
  • State-of-the-art performance and scalability. Through experiments, we demonstrate that UM outperforms both traditional optimization (by 24–82%) and single-task DRL methods (by 6–14%) in terms of solution quality while maintaining real-time inference (1–10 seconds) for networks with up to 1,000 nodes. Sensitivity analyses confirm the model’s robustness under varying operational conditions, and validation on the real-world road network underscores the model’s practical scalability.
@@ 2 Related work
PDRA has evolved with technological advancements. Traditional assessment methods, such as field surveys (Arii 2013) and satellite remote sensing (Bravo et al. 2019), are constrained by slow deployment, labor intensity, and low resolution, which hinder timely decision-making in crisis scenarios. To overcome these limitations, drones have emerged as a transformative solution for PDRA. Drones offer strong mobility, cost efficiency, and high-resolution imaging capabilities, enabling rapid evaluation of road damage (e.g., cracks or landslides) and operability, thus becoming a cornerstone of modern PDRA (Enayati et al. 2023, Zhang et al. 2023).
A critical challenge in PDRA is solving drone routing for large-scale road networks. Existing approaches, including exact algorithms and metaheuristics, face two key limitations: computational inefficiency with increasing problem scale and heavy reliance on domain expertise in model and solution approach design (Gong et al. 2025, Jiang et al. 2025). For instance, Zhang et al. (2023) developed a branch-and-price model validated on 107-node road networks, but its solution time grows exponentially as scale increases. Similarly, Adsanver et al. (2025)’s heuristic approach, though effective for 148-node road networks, requires extensive domain knowledge and scales poorly for real-world large-scale PDRA. To address these issues, Gong et al. (2025) introduced a DRL framework for PDRA, leveraging an attention-based encoder-decoder model (AEDM). This method eliminates domain knowledge dependencies by learning routing strategies directly from data, achieving real-time inference (1–10 seconds) for networks with up to 1,000 nodes and outperforming commercial solvers by 16–69%.
Like vehicle routing problems with diverse variants (Luo et al. 2023, Zhou et al. 2024, Liu et al. 2024, Berto et al. 2024), PDRA involves multiple routing variants shaped by operational constraints: open routes (no depot return), time windows (temporal assessment constraints), and multi-depot scenarios (distributed launch points). However, existing methods, including exact algorithms, heuristics, and AEDM proposed by Gong et al. (2025), operate in a single-task paradigm, requiring separate models for each variant. Adapting to new constraints (e.g., switching from closed to open routes) necessitates full model redesign, leading to inefficiencies and limited adaptability to emerging PDRA scenarios. Table 1 compares these methods, highlighting gaps in unified modeling and adaptability. These approaches lack the ability to handle multiple variants or unseen attributes, underscoring the need for flexible frameworks that can address such limitations.
@@
Oruc and Kara (2018) Max. value & profit Base route heuristic 44 Slow Yes
Glock and Meyer (2020) Max. info. in time Neighborhood search 625 Slow Yes
Zhang et al. (2023) Max. reward Branch-and-price 107 Slow Yes
Yin et al. (2023) Min. cost Branch-and-price 45 Slow Yes
Morandi et al. (2024) Max. prize Branch-and-cut 50 Slow Yes
Adsanver et al. (2025) Max. priority Variable neighborhood 148 Slow Yes
Gong et al. (2025) Max. reward DRL (AEDM) 1,000 Fast No
@@ 3 Preliminaries
In this section, we briefly introduce the PDRA using drones. For detailed elaboration, please refer to Gong et al. (2025). To maintain consistency and facilitate easy reference, we adopt their notation as much as possible. We first present the drone routing problem formulation, followed by the network transformation method and instance generation procedure.
@@ 3.1 Drone routing
Each drone is equipped with high-definition cameras and operates under three critical constraints for PDRA. First, each drone has a battery-imposed flight time limit $Q$. Second, the entire assessment operation must be completed within a maximum allowable time $p_{\text{max}}$. Third, due to operational altitude and visual range limitations, a bidirectional road link can only be assessed if a drone flies directly along it (Zhang et al. 2023). The PDRA problem involves two distinct time considerations: the computational time to generate assessment routes and the operational time for drones to execute those routes.
@@ 3.2 Network transformation
To resolve the ambiguities in the dual-network formulation, where both road links and direct routes may connect the same pair of nodes, Gong et al. (2025) developed a network transformation method tailored for PDRA, converting the link-based routing problem into an equivalent node-based formulation. This transformation, illustrated in Figure 1(b), involves splitting each road link $(i,j)\in A$ by introducing an artificial node $p\in P$, creating two new links: $(i,p)$ and $(p,j)$. The transformation process converts the dual network $G^{\prime\prime}=(N,A+A^{\prime})$ into a reduced network $\bar{G}=(\bar{N},\bar{A})=(N\cup P,2A\cup A^{\prime})$, where $|P|=|A|$. In the transformed network, the ambiguity of multiple edges between nodes is resolved to support accurate PDRA: traversing $(i,p)$ and $(p,j)$ corresponds to assessing the original link $(i,j)$, while direct links in $A^{\prime}$ remain unchanged for transit purposes. As shown in Figure 1(c) and (d), the information value of artificial node $p$ is set as $c_{p}=c_{ij}$, and the traversal time is split as $t_{ip}=t_{pj}=t_{ij}/2$, effectively converting the link-based PDRA routing problem into a node-based problem.
This transformation allows the PDRA problem to be formulated as a variant of the orienteering problem where artificial nodes in $P$ carry information values (prizes) and must be visited to collect damage assessment data, while original nodes in $N$ serve as waypoints for routing. However, critical differences exist from standard orienteering problems to accommodate PDRA needs: (1) original nodes in $N$ carry no information value, (2) artificial nodes in $P$ are isolated from each other and connect only to specific nodes in $N$, and (3) artificial nodes in $P$ are exclusive (each can be visited by only one drone) to avoid redundant assessments, whereas nodes in $N$ can be revisited, enabling loop formations in drone paths to enhance PDRA efficiency.
@@ 3.3 Instance generation
Because large-scale real-world road network datasets suitable for training deep learning models are scarce, Gong et al. (2025) developed a synthetic road network generation method specifically tailored for drone routing problems in PDRA. The generation process follows a four-stage pipeline: (1) Grid Network Initialization, which creates a uniform square grid of $N$ nodes over the coordinate space $0,1]^{2}$; (2) Link Pruning, which randomly removes edge subsets while maintaining network connectivity to simulate real-world road networks; (3) Node Perturbation, which introduces bounded random perturbations to node coordinates to avoid regularity, ensuring the synthetic networks better reflect the irregularities of actual road networks; and (4) Attribute Assignment, which computes Euclidean lengths for remaining edges and assigns random damage information values to transformed artificial nodes.
@@ 4 Variants and motivations
Drone routing for PDRA encompasses diverse problem variants, each reflecting unique operational constraints or environmental realities in post-disaster scenarios. These variants as listed below extend the preliminary PDRA (Gong et al. 2025) and shape the need for a UM to enhance practical applicability.
@@
The above variants create distinct problem structures for PDRA. The presence or absence of the three attributes results in 8 variants as shown in Table 2. Existing methods, including exact methods, heuristics, and AEDM proposed by Gong et al. (2025), operate in a single-task paradigm, requiring separate models for each PDRA variant, which hinders practical deployment. A UM capable of solving all PDRA variants within a single model addresses these shortcomings by reducing computational overhead and improving deployment flexibility. Moreover, operational environments may introduce future, previously unseen attributes in evolving PDRA scenarios. In this study, MD is deliberately regarded as such a future attribute for PDRA. After training UM (discussed in Section 5), the attribute can be incorporated via a lightweight adapter layer mechanism (discussed in Section 6), enabling efficient finetuning without full retraining. This design ensures adaptability to evolving post-disaster response requirements, a critical feature for robust PDRA systems.

### v2

@@
Abstract
Post-disaster road assessment (PDRA) is essential for emergency response, enabling rapid evaluation of infrastructure conditions and efficient allocation of resources. Although drones provide a flexible and effective tool for PDRA, routing them in large-scale networks remains challenging. Exact and heuristic optimization methods scale poorly and demand domain expertise, while existing deep reinforcement learning (DRL) approaches adopt a single-task paradigm, requiring separate models for each problem variant and lacking adaptability to evolving operational needs. This study proposes a unified model (UM) for drone routing that simultaneously addresses eight PDRA variants. By training a single neural network across multiple problem configurations, UM captures shared structural knowledge while adapting to variant-specific constraints through a modern transformer encoder-decoder architecture. A lightweight adapter mechanism further enables efficient finetuning to unseen attributes without retraining, enhancing deployment flexibility in dynamic disaster scenarios. Extensive experiments demonstrate that the UM reduces training time and parameters by a factor of eight compared with training separate models, while consistently outperforming single-task DRL methods by 6–14%, heuristic algorithms by 22–42%, and commercial solvers by 24–82% in terms of solution quality (total collected information value). The model achieves rapid solutions (1–10 seconds) across networks of up to 1,000 nodes, with robustness confirmed through sensitivity analyses. Moreover, finetuning experiments show that unseen attributes can be effectively incorporated with minimal cost while retaining high solution quality. Validation on a real-world transportation network further demonstrates the model’s practical scalability. The proposed UM advances neural combinatorial optimization for time-critical applications, offering a computationally efficient, high-quality, and adaptable solution for drone-based PDRA. The source code for UM is publicly available at https://github.com/PJ-HTU/UM_PDRA.
keywords:
Post-disaster road assessment; Drone routing; Deep reinforcement learning; Multi-task learning; Lightweight finetuning; Neural combinatorial optimization
@@ 1 Introduction
Disasters, such as earthquakes and hurricanes, pose severe threats to human lives, infrastructure, and economic stability (avishan2023humanitarian). In the immediate aftermath of such events, uncertainties regarding transportation infrastructure disruptions emerge, encompassing the extent of road network damage, operability, and feasibility of short-term restoration. Therefore, rapid assessment of road network damage is critical for effective emergency response, as it enables timely delivery of supplies, evacuation planning, and resource allocation (akbari2021online; zhang2023robust). This process, referred to as post-disaster road assessment (PDRA), is hindered by traditional ground-based methods, which are often impeded by impassable roads, hazardous conditions, and delays, making them impractical for time-sensitive disaster scenarios (zhang2023robust; Gong2025Deep). Unmanned aerial vehicles (or drones) have emerged as a transformative solution for PDRA, offering high mobility, rapid deployment, and the ability to capture high-resolution imagery in hazardous environments (hong2018range; pang2022adaptive; van2025stochastic). Equipped with advanced sensors, drones can efficiently assess road damage severity, operability, and repair feasibility, providing critical data to guide relief efforts. Examples include their use in post-typhoon Haiyan (ezequiel2015aerial) and the Nepal earthquake (abuali2025innovative).
@@
Despite these advantages, solving drone routing for large-scale PDRA remains challenging. Traditional approaches, including exact methods (e.g., branch-and-bound algorithms) (huang2013continuous; zhang2023robust; yin2023robust; morandi2024orienteering) and heuristic methods (e.g., simulated annealing) (oruc2018post; glock2020mission; adsanver2025predictive), suffer from significant limitations in real-world large-scale PDRA. These methods exhibit solution times that grow dramatically with problem complexity, making them ill-suited to time-critical disaster response where delays can be life-threatening. Moreover, developing effective solving algorithms requires extensive domain knowledge and substantial time investment in algorithm design. To address these two issues, Gong2025Deep proposed a deep reinforcement learning (DRL) approach for PDRA, which learns high-quality routing strategies directly from data without relying on domain knowledge, achieving rapid inference (1–10 seconds) and outperforming commercial solvers in solution quality. In recent years, DRL has been used to learn neural network-based heuristics for solving combinatorial optimization problems, receiving growing research attention for its potential to generate high-quality solutions with minimal human effort (vinyals2015pointer; bello2016neural; kool2018attention; kwon2020pomo; luo2023neural; zhou2024mvmoe; liu2024multi; berto2024routefinder). However, this DRL method operates in a single-task paradigm, requiring a separate model for each specific PDRA variant. When problem characteristics change, such as variations in objective functions or constraints, a new model should be trained from scratch, leading to high computational costs and deployment inefficiencies. Furthermore, the existing method lacks adaptability to unseen future attributes in evolving PDRA scenarios.
@@
Inspired by advancements in large language models (LLMs), where a single model can be applied to diverse tasks with remarkable adaptability and deployment efficiency (achiam2023gpt; touvron2023llama; guo2025deepseek; huang2025orlm; ling2026review), this study extends the philosophy to neural combinatorial optimization for PDRA. The main contributions are summarized as follows:
@@
  • State-of-the-art performance and scalability. Through extensive experiments, we demonstrate that UM outperforms single-task DRL methods (by 6–14%), heuristic algorithms (by 22–42%), and commercial solvers (by 24–82%) in terms of solution quality while maintaining rapid inference (1–10 seconds) for networks with up to 1,000 nodes. Sensitivity analyses confirm the model’s robustness under varying operational conditions, and validation on the real-world road network underscores the model’s practical scalability.
@@ 2 Related work
PDRA has evolved with technological advancements. Traditional assessment methods, such as field surveys (maya2013rapid) and satellite remote sensing (bravo2019use), are constrained by slow deployment, labor intensity, and low resolution, which hinder timely decision-making in crisis scenarios. To overcome these limitations, drones have emerged as a transformative solution for PDRA. Drones offer strong mobility, cost efficiency, and high-resolution imaging capabilities, enabling rapid evaluation of road damage (e.g., cracks or landslides) and operability, thus becoming a cornerstone of modern PDRA (hong2018range; zhang2023robust; bogyrbayeva2023deep).
A critical challenge in PDRA is solving drone routing for large-scale road networks. Existing approaches, including exact algorithms and metaheuristics, face two key limitations: computational inefficiency with increasing problem scale and heavy reliance on domain expertise in model and solution approach design (Gong2025Deep; jiang2025large). For instance, zhang2023robust developed a branch-and-price model validated on 107-node road networks, but its solution time grows exponentially as scale increases. Similarly, adsanver2025predictive’s heuristic approach, though effective for 148-node road networks, requires extensive domain knowledge and scales poorly for real-world large-scale PDRA. To address these issues, Gong2025Deep introduced a DRL framework for PDRA, leveraging an attention-based encoder-decoder model (AEDM). This method eliminates domain knowledge dependencies by learning routing strategies directly from data, achieving rapid inference (1–10 seconds) for networks with up to 1,000 nodes and outperforming commercial solvers by 16–69%.
Like vehicle routing problems with diverse variants (luo2023neural; zhou2024mvmoe; liu2024multi; berto2024routefinder), PDRA involves multiple routing variants shaped by operational constraints: open routes (no depot return), time windows (temporal assessment constraints), and multi-depot scenarios (distributed launch points). However, existing methods, including exact algorithms, heuristics, and AEDM proposed by Gong2025Deep, operate in a single-task paradigm, requiring separate models for each variant. Adapting to new constraints (e.g., switching from closed to open routes) necessitates full model redesign, leading to inefficiencies and limited adaptability to emerging PDRA scenarios. Table 1 compares these methods, highlighting gaps in unified modeling and adaptability. These approaches lack the ability to handle multiple variants or unseen attributes, underscoring the need for flexible frameworks that can address such limitations.
@@
oruc2018post Max. value & profit Base route heuristic 44 Slow Yes
glock2020mission Max. info. in time Neighborhood search 625 Slow Yes
zhang2023robust Max. reward Branch-and-price 107 Slow Yes
yin2023robust Min. cost Branch-and-price 45 Slow Yes
morandi2024orienteering Max. prize Branch-and-cut 50 Slow Yes
adsanver2025predictive Max. priority Variable neighborhood 148 Slow Yes
Gong2025Deep Max. reward DRL (AEDM) 1,000 Fast No
@@ 3 Preliminaries
In this section, we briefly introduce the PDRA using drones. For detailed elaboration, please refer to Gong2025Deep. To maintain consistency and facilitate easy reference, we adopt their notation as much as possible. We first present the drone routing problem formulation, followed by the network transformation method and instance generation procedure.
@@ 3.1 Drone routing
Each drone is equipped with high-definition cameras and operates under three critical constraints for PDRA. First, each drone has a battery-imposed flight time limit $Q$. Second, the entire assessment operation must be completed within a maximum allowable time $p_{\text{max}}$. Third, due to operational altitude and visual range limitations, a bidirectional road link can only be assessed if a drone flies directly along it (zhang2023robust). The PDRA problem involves two distinct time considerations: the computational time to generate assessment routes and the operational time for drones to execute those routes.
@@ 3.2 Network transformation
To resolve the ambiguities in the dual-network formulation, where both road links and direct routes may connect the same pair of nodes, Gong2025Deep developed a network transformation method tailored for PDRA, converting the link-based routing problem into an equivalent node-based formulation. This transformation, illustrated in Figure 1 (b), involves splitting each road link $(i,j)\in A$ by introducing an artificial node $p\in P$, creating two new links: $(i,p)$ and $(p,j)$. The transformation process converts the dual network $G^{\prime\prime}=(N,A+A^{\prime})$ into a reduced network $\bar{G}=(\bar{N},\bar{A})=(N\cup P,2A\cup A^{\prime})$, where $|P|=|A|$. In the transformed network, the ambiguity of multiple edges between nodes is resolved to support accurate PDRA: traversing $(i,p)$ and $(p,j)$ corresponds to assessing the original link $(i,j)$, while direct links in $A^{\prime}$ remain unchanged for transit purposes. As shown in Figure 1 (c) and (d), the information value of artificial node $p$ is set as $c_{p}=c_{ij}$, and the traversal time is split as $t_{ip}=t_{pj}=t_{ij}/2$, effectively converting the link-based PDRA routing problem into a node-based problem.
This transformation might allow the PDRA problem to be formulated as a variant of the orienteering problem (OP) (kobeaga2018efficient) where artificial nodes in $P$ carry information values (prizes) and must be visited to collect damage assessment data, while original nodes in $N$ serve as waypoints for routing. However, critical differences exist from standard OP to accommodate PDRA needs: (1) original nodes in $N$ carry no information value, (2) artificial nodes in $P$ are isolated from each other and connect only to specific nodes in $N$, and (3) artificial nodes in $P$ are exclusive (each can be visited by only one drone) to avoid redundant assessments, whereas nodes in $N$ can be revisited, enabling loop formations in drone paths to enhance PDRA efficiency. This characteristic differs from OP’s Miller-Tucker-Zemlin (MTZ) formulation (miller1960integer), which eliminates loops to maintain route connectivity. These architectural distinctions indicate that existing OP methods (whether exact or approximate) cannot be readily adapted. This underscores a prevalent challenge with conventional algorithmic approaches. Even modest alterations to problem characteristics frequently necessitate extensive reformulation, demanding considerable effort and deep domain expertise (luo2023neural; zhou2024mvmoe; liu2024multi; berto2024routefinder; yang2025heuragenix; liu2025eoh).
@@ 3.3 Instance generation
Because large-scale real-world road network datasets suitable for training deep learning models are scarce, Gong2025Deep developed a synthetic road network generation method specifically tailored for drone routing problems in PDRA. The generation process follows a four-stage pipeline: (1) Grid Network Initialization, which creates a uniform square grid of $N$ nodes over the coordinate space $0,1]^{2}$; (2) Link Pruning, which randomly removes edge subsets while maintaining network connectivity to simulate real-world road networks; (3) Node Perturbation, which introduces bounded random perturbations to node coordinates to avoid regularity, ensuring the synthetic networks better reflect the irregularities of actual road networks; and (4) Attribute Assignment, which computes Euclidean lengths for remaining edges and assigns random damage information values to transformed artificial nodes.
@@ 4 Variants and motivations
Drone routing for PDRA encompasses diverse problem variants, each reflecting unique operational constraints or environmental realities in post-disaster scenarios. These variants as listed below extend the preliminary PDRA (Gong2025Deep) and shape the need for a UM to enhance practical applicability.
@@
The above variants create distinct problem structures for PDRA. The presence or absence of the three attributes results in 8 variants as shown in Table 2. Existing methods, including exact methods, heuristics, and AEDM proposed by Gong2025Deep, operate in a single-task paradigm, requiring separate models for each PDRA variant, which hinders practical deployment. A UM capable of solving all PDRA variants within a single model addresses these shortcomings by reducing computational overhead and improving deployment flexibility. Moreover, operational environments may introduce future, previously unseen attributes in evolving PDRA scenarios. In this study, MD is deliberately regarded as such a future attribute for PDRA. After training UM (discussed in Section 5), the attribute can be incorporated via a lightweight adapter layer mechanism (discussed in Section 6), enabling efficient finetuning without full retraining. This design ensures adaptability to evolving post-disaster response requirements, a critical feature for robust PDRA systems.